2) You are the manager of a firm in a new industry. You have gotten the jump on the only other producer in the market. You kn...
Answer #1

= 1250-P/2 P= (1250-0) & P= 2500-2012 Q = 4 + Q 2, - de MOA = 8 MC2 = 6. Stackelberg Game 1 (A) moves Tg = (2500-29-292-6) 927,= ( 1245-24 + 2) = ( 1245-49 -0. anilan 1245-29 U= 629.5 qg= 6 23.5 - .5(692.5) - 318.25 80 8= 4+Q2 = 934.75 P= 2 500 - 2